# Latent Intrinsics Emerge from Training to Relight

### v1

@@ 1 Introduction
One common strategy to relight a scene is to infer scene characteristics (geometry, surface properties) using inverse graphics methods, then render the scene with a new light source. This approach is fraught with difficulties, including the challenge of selecting which material properties to infer and managing error propagation. These methods perform best in outdoor scenes with significant shadow movements but struggle with indoor scenes where interreflections create complex effects (Section 4.2).
As this paper demonstrates, a purely data-driven method offers an attractive alternative. A source scene (represented by an image) is encoded to produce a latent representation of intrinsic scene properties. A source illumination (represented by another image) is encoded to produce a latent representation of illumination properties. These intrinsic and extrinsic properties are combined and then decoded to produce the relighted image. As a byproduct of this training, we find that the latent representation of intrinsic scene properties behaves like an albedo, while another latent representation acts as a lighting controller.
@@ 2 Related Work
Intrinsic Images. Humans have been known to perceive scene properties independent of lighting since at least 1867 [45, 20, 4, 19]. In computer vision, the idea dates to Barrow and Tenenbaum [3] and comprises at least depth, normal, albedo, and surface material maps. Depth and normal estimation are now well established (eg [23]). There is a rich literature on albedo estimation (dating to 1959 [29, 30]!). A detailed review appears in [15], which breaks out methods as to what kinds of training data they see. Early methods do not see any form of training data, but more recently both CGI data and manual annotations of relative lightness (labels) have become available. Early efforts, such SIRFS [1], focused on using shading information to recover shape, illumination, and reflectance, highlighting the importance of modeling these factors for intrinsic image analysis. Recent strategies include: deep networks trained on synthetic data [32, 22, 14]; and conditional generative models [27].
@@
Using Intrinsic Images for Relighting. Bhattad and Forsyth [6] demonstrated that intrinsic images could be used for reshading inserted objects. This approach can be extended by adjusting the shading in both the foreground and background to eliminate discrepancies [11]. Intrinsic images and geometry-aware networks have been used for multi-view relighting [40]. StyLitGAN [8] introduced a method to relight images by identifying directional vectors in the latent space of StyleGAN, but can only relight StyleGAN generated images and requires explicit albedo and shading to guide relighting. LightIt [28] controls lighting changes in image generation using diffusion models, by conditioning on shading and normal maps to achieve consistent and controllable lighting. Like these methods, we use intrinsics and extrinsics to relight, but ours are latent, with no explicit physical meaning.
@@
Figure 2: The network diagram of our relighting model. The model functions as an autoencoder, comprising an encoder ${\bm{E}}$ and a decoder ${\bm{D}}$. Left Half: The encoder ${\bm{E}}$ maps input image ${\bm{I}}_{s}^{l}$, captured under scene $s$ and lighting $l$, to low-dimensional extrinsic features ${\bm{L}}_{s}^{l}$ and set of intrinsic features map $\{S_{s,i}^{l}\}_{i}$. The decoder ${\bm{D}}$ then generates new images based on these intrinsic and extrinsic representations. Right Half: We employ constrained scaling for the injection of ${\bm{L}}_{s}^{l}$, utilizing $0<\alpha\ll 1$ to regularize the information passed from ${\bm{L}}_{s}^{l}$, thereby enforcing a low-dimensional parameterization of the extrinsic features. We train our system to relight target images given input paired of image captured under the same scene $s$. During inference, our system demonstrates the capability to generalize to arbitrary reference images for relighting and can estimate albedo for free.
@@ 3.1 Model structure
We apply L2 normalization along the feature channel to both sets of features. During training, we introduce random Gaussian noise to the input image to enhance semantic scene understanding capabilities:
@@ 4.1 Experiment Details
Training Details We train our model using the MIT multi-illumination dataset [37], which includes images of 1,015 indoor scenes captured under 25 fixed lighting patterns, totaling 25,375 images. We follow the official data split and train our model on the 985 training scenes. During training, we randomly sample pairs of images from the same scene under different lighting conditions and perform random spatial cropping, with the crop ratio randomly selected between 0.2 and 1.0, followed by resizing the cropped image to a resolution of 256x256. For further details, please refer to our appendix.
@@ 4.2 Evaluating image relighting
Figure 3: Our method surpasses all other approaches in estimating light and rendering the scene. The Unsupervised SA-AE [21] method fails by incorporating intrinsic elements from reference images. The S3Net [50] approach struggles with rendering when using unpaired reference images. Right: A zoomed-in view of the chrome ball used as a probe to evaluate detail preservation in the environment map. Our method effectively retains the intricate room layout and accurately renders the appropriate lighting patterns.
@@
Figure 5: Qualitative results for relighting interior scenes using our relighter trained on images obtained from StyLitGAN (which produces multiple illuminations of a generated scene). StyLitGAN has a strong tendency to increase or decrease illumination by adjusting luminaires, typically bedside lights but also light coming through french windows, etc. On the left, where the reference lighting tends to be brighter and more concentrated, notice how for the two top images, our relighter has identified and "turned up" the bedside lights; for the third, it has resisted StyLitGAN’s tendency to invent helpful luminaires (there isn’t a bedside light where StyLitGAN imputed one, as close inspection shows). On the right, where the reference lighting is much more uniform, our relighter has achieved this by "turning down" bedside lights. This is an emergent phenomenon; the method is not supplied with any explicit luminaire model, etc.
@@
tableWe benchmark our albedo esimation on test set of IIW dataset [5] and compare with others, though the reliability has been questioned by recent papers [15]. Flat denotes postprocessing images with flattening [9]. Despite our model never being trained on albedo maps or CG data, our best configuration significantly outperforms all other methods suggesting our model learns high quality intrinsic representations
@@
Figure 6: Qualitative Comparison of Albedo Estimation on the IIW Dataset. Although our model has never been trained on any albedo-like maps, it effectively removes the effects of external light and dark shadows from the input. In contrast, Intrinsic Diffusion [27], a supervised method trained on computer graphics data, often produces color-drifted estimations, likely due to the domain shift between CG data and real images.
@@ 4.3 Zero-shot albedo evaluation
The resulting classification (one lighter than two; two lighter than one; equivalent) is then compared to human annotations $J$ using the confidence score $w_{i}$ for each annotation pair. We report WHDR on the IIW test split in Table 4.2 to facilitate comparison with other approaches. Since our model is not trained with any albedo maps or computer-generated images, we need to adjust the threshold for the optimal performance. Following prior work, we optimize $\delta$ on the training split, which significantly improves our performance from 28.97 to 19.09. Additionally, we enhance our performance by post-processing our albedo map using flattening [9], an optimization technique to further reduce color variations. With this improvement, our results reach 15.81, substantially outperforming the intrinsic diffusion model  [27], a diffusion-based albedo regression model trained on computer graphics data. In Figure 6, we show some qualitative comparisons to intrinsic diffusion. We observe that our method effectively removes external lighting effects and does not suffer from color-drift due to domain gap unlike intrinsic diffusion, which is trained on CG data.
@@ 5 Discussion, Limitations and Future Work
Our method presents a significant advancement in image relighting by demonstrating that intrinsic properties such as albedo can emerge naturally from training on relighting tasks without explicit supervision. This finding simplifies the relighting process, eliminating the need for detailed geometric and surface models and enhancing the model’s ability to generalize across diverse and unseen scenes. By encoding scene and illumination properties as latent representations, we achieve accurate and flexible relighting, applicable to various fields such as virtual reality and cinematic post-production. This approach reduces the learning process’s complexity and offers a new perspective on designing deep learning models to capture and utilize intrinsic scene properties. These findings can guide future research toward more efficient and scalable relighting techniques, encouraging the development of models that can handle various lighting conditions and scene complexities.
@@
Nonetheless, there is good evidence that explicit intrinsic information can be extracted from our latent intrinsics. Our method clearly “knows” albedo, and this information can be elicited without examples. Similarly, it “knows” something about luminaires, such as their locations and effects. It is intriguing to speculate that it “knows” other information relevant to relighting, such as depth or surface microstructure. Future work will pursue this line of inquiry further and also focus on developing a purely unsupervised framework to infer intrinsic and extrinsic properties from collections of in-the-wild images. This will include refining probing techniques for better extraction of explicit intrinsics and identifying additional intrinsic properties crucial for relighting that do not align with the current taxonomy. We believe this will improve the applicability and robustness of our approach, making it suitable for a wider range of real-world scenarios.
@@ Acknowledgment
AB thanks Stephan R. Richter for the discussions that led to the consideration of intrinsic images as latent variables.

### v2

@@ 1 Introduction
One common approach to relighting a scene is to infer scene characteristics (geometry, surface properties) using inverse graphics methods, then render the scene with a new light source. This approach is fraught with difficulties, including the challenge of selecting which material properties to infer and managing error propagation. These methods perform best in outdoor scenes with significant shadow movements but struggle with indoor scenes where interreflections create complex effects (Section 4.2).
As this paper demonstrates, a purely data-driven method offers an attractive alternative. A source scene, represented by an image, is encoded to produce a latent representation of intrinsic scene properties. A source illumination, represented by another image, is encoded to produce a latent representation of illumination properties. These intrinsic and extrinsic properties are combined and then decoded to produce the relighted image. As a byproduct of this training, we find that the latent representation of intrinsic scene properties behaves like an albedo, while another latent representation acts as a lighting controller.
@@ 2 Related Work
Intrinsic Images. Humans have been known to perceive scene properties independent of lighting since at least 1867 [46, 21, 4, 20]. In computer vision, the idea dates to Barrow and Tenenbaum [3] and comprises at least depth, normal, albedo, and surface material maps. Depth and normal estimation are now well established (eg [24]). There is a rich literature on albedo estimation (dating to 1959 [30, 31]!). A detailed review appears in [16], which breaks out methods as to what kinds of training data they see. Early methods do not see any form of training data, but more recently both CGI data and manual annotations of relative lightness (labels) have become available. Early efforts, such as SIRFS [1], focused on using shading information to recover shape, illumination, and reflectance, highlighting the importance of modeling these factors for intrinsic image analysis. Recent strategies include: deep networks trained on synthetic data [33, 23, 15]; and conditional generative models [29].
@@
Using Intrinsic Images for Relighting. Bhattad and Forsyth [6] demonstrated that intrinsic images could be used for reshading inserted objects. This approach can be extended by adjusting the shading in both the foreground and background to eliminate discrepancies [12]. Intrinsic images and geometry-aware networks have been used for multi-view relighting [41]. StyLitGAN [9] introduced a method to relight images by identifying directional vectors in the latent space of StyleGAN, but can only relight StyleGAN generated images and requires explicit albedo and shading to guide relighting. It can be extended to real images using a GAN inversion, but does not generalize [7]. LightIt [28] controls lighting changes in image generation using diffusion models, by conditioning on shading and normal maps to achieve consistent and controllable lighting. Like these methods, we use intrinsics and extrinsics to relight, but ours are latent, with no explicit physical meaning.
@@
Figure 2: The network diagram of our relighting model. The model functions as an autoencoder, comprising an encoder ${\bm{E}}$ and a decoder ${\bm{D}}$. Left Half: The encoder ${\bm{E}}$ maps input image ${\bm{I}}_{s}^{l}$, captured under scene $s$ and lighting $l$, to low-dimensional extrinsic features ${\bm{L}}_{s}^{l}$ and set of intrinsic features map $\{S_{s,i}^{l}\}_{i}$. The decoder ${\bm{D}}$ then generates new images based on these intrinsic and extrinsic representations. Right Half: We employ constrained scaling for the injection of ${\bm{L}}_{s}^{l}$, utilizing $0<\alpha\ll 1$ to regularize the information passed from ${\bm{L}}_{s}^{l}$, thereby enforcing a low-dimensional parameterization of the extrinsic features. We train our system to relight target images given input paired with images captured under the same scene $s$. During inference, our model demonstrates the ability to generalize to arbitrary reference images for relighting and can estimate albedo for free.
@@ 3.1 Model structure
We apply L2 normalization along the feature channel to both sets of features. During training, we add random Gaussian noise to the input image to enhance semantic scene understanding capabilities:
@@ 4 Experiments
We will first provide a brief description of our experimental procedure (Sec 4.1), followed by a discussion on how we evaluate the various relighting capabilities of our approach, including its strong generalization across datasets with different distributions (Sec 4.2). Finally, we will present the emergent albedo that is recovered from the latent intrinsic without using any albedo-like images (Sec 4.3).
@@ 4.1 Experiment Details
Training Details We train our model using the MIT multi-illumination dataset [38], which includes images of 1,015 indoor scenes captured under 25 fixed lighting, totaling 25,375 images. We follow the official data split and train our model on the 985 training scenes. During training, we randomly sample pairs of images from the same scene under different lighting conditions and perform random spatial cropping, with the crop ratio randomly selected between 0.2 and 1.0, followed by resizing the cropped image to a resolution of 256x256. For further details, please refer to our appendix.
@@ 4.2 Evaluating image relighting
Figure 3: Our method outperforms all other approaches in estimating light and rendering the scene. The Unsupervised SA-AE [22] method fails by incorporating intrinsic elements from reference images. The S3Net [51] approach struggles with rendering when using unpaired reference images. Right: A zoomed-in view of the chrome ball was used as a probe to evaluate detail preservation in the environment map. Our method effectively retains the intricate room layout and accurately renders the appropriate lighting effects.
@@
Figure 5: Qualitative results for relighting interior scenes using our relighter trained on images obtained from StyLitGAN (which produces multiple illuminations of a generated scene). StyLitGAN has a strong tendency to increase or decrease illumination by adjusting luminaires, typically bedside lights but also light coming through French windows, etc. On the left, where the reference lighting tends to be brighter and more concentrated, notice how for the two top images, our relighter has identified and "turned up" the bedside lights; for the third, it has resisted StyLitGAN’s tendency to invent helpful luminaires (there isn’t a bedside light where StyLitGAN imputed one, as close inspection shows). On the right, where the reference lighting is much more uniform, our relighter has achieved this by "turning down" bedside lights. This is an emergent phenomenon; the method is not supplied with any explicit luminaire model or labeled data.
@@
Input   Refer to caption
Ref   Refer to caption
Relight   Refer to caption
Input   Refer to caption
Ref   Refer to caption
Relight   Refer to caption
Input   Refer to caption
Ref   Refer to caption
Relight   Refer to caption
Figure 6: Zero-Shot Relighting. Our relighting model, trained only on the multi-illumination dataset, generalizes well to out-of-distribution images, as shown on the IIW dataset (first row) and StyleGAN images (second row). It accurately infers scene geometry and lighting. Note that it identifies and turns on the bedside lamps in StyleGAN images despite having no training in bedroom images. This demonstrates the model’s strong generalization ability and the model clearly “knows" something about light sources.
Zero-Shot Relighting: In Figure.6, we show our model’s strong generalization by applying the model solely trained on multi-illumination dataset—without additional training or fine-tuning—to relight IIW and StyleGAN-generated images. Despite the significant distribution shift in lighting patterns and room setup, our model accurately identifies luminaires and relights images.
@@
tableWe benchmark our albedo esimation on test set of IIW dataset [5] and compare with others, though the reliability has been questioned by recent papers [16]. Flat denotes postprocessing images with flattening [10]. Despite our model never being trained on albedo maps or CG data, our best configuration significantly outperforms all other methods suggesting our model learns high-quality intrinsic representations
@@
Figure 7: Qualitative Comparison of Emergent Albedo from Latent Intrinsics on the IIW Dataset. Although our model has never been trained on any albedo-like maps, it effectively removes the effects of external light and dark shadows from the input. In contrast, Intrinsic Diffusion [29], a supervised method trained on large computer graphics data, often produces color-drifted estimations, likely due to the domain shift between CG data and real images. Observe the subdued lighting around the mirrors (top row, right) in our recovered albedo. Also, pay attention to all the details inside the refrigerator, which are visible in our recovered albedos (bottom row; right) compared to intrinsic diffusion. For comparison, we also display naive flattening (in the second column), which by itself cannot effectively reduce the strong lighting effects.
@@ 4.3 Zero-shot albedo evaluation
The resulting classification (one lighter than two; two lighter than one; equivalent) is then compared to human annotations $J$ using the confidence score $w_{i}$ for each annotation pair. We report WHDR on the IIW test split in Table 4.2 to facilitate comparison with other approaches. Since our model is not trained with any albedo maps or computer-generated images, we need to adjust the threshold for the optimal performance. Following prior work, we optimize $\delta$ on the training split, which significantly improves our performance from 28.97 to 19.09. Additionally, we enhance our performance by post-processing our albedo map using flattening [10], an optimization technique to further reduce color variations. With this improvement, our results reach 15.81, substantially outperforming the intrinsic diffusion model  [29], a diffusion-based albedo regression model trained on computer graphics data. In Figure 7, we show some qualitative comparisons to intrinsic diffusion. We observe that our method effectively removes external lighting effects and does not suffer from color drift due to domain gap unlike intrinsic diffusion, which is trained on CG data.
Ours        Intrinsic        Image
Diffusion
Figure 8: Qualitative comparison of albedo stability under varying lighting conditions. Images shown are from the multi-illumination dataset test split. The top row features images under different lighting environments. The middle row presents estimated albedos obtained from Intrinsic Diffusion [29], while the bottom row shows the recovered albedos from the latent intrinsic representation. Intrinsic Diffusion has large color drift and is sensitive to changes in lighting. In contrast, the albedos recovered from latent intrinsics remain stable under lighting changes, even in extreme conditions.
Sensitivity to light changes: Albedo are scene properties that are independent of lighting changes. In Figure. 8, we qualitatively assess this characteristic by varying lighting conditions, comparing our approach with the state-of-the-art supervised method, Intrinsic Diffusion [29]. Our method demonstrates consistent and accurate estimations that remain stable even under extreme lighting variations. In contrast, Intrinsic Diffusion [29] shows significant deviation from the natural color distribution and are sensitive to lighting changes.
@@ 5 Discussion, Limitations and Future Work
Our method presents an important advancement in image relighting by demonstrating that intrinsic properties such as albedo can emerge naturally from training on relighting tasks without explicit supervision. This finding simplifies the relighting process, eliminating the need for detailed geometric and surface models and enhancing the model’s ability to generalize across diverse and unseen scenes. By encoding scene and illumination properties as latent variables, we achieve accurate and flexible relighting. Our findings will have implications for various fields such as virtual reality and cinematic post-production. This approach reduces the learning process’s complexity and offers a new perspective on designing deep learning models to capture and utilize intrinsic scene properties. These findings can guide future research toward a more efficient and scalable relighting approach, encouraging the development of models that can handle various lighting conditions and scene complexities.
@@
Nonetheless, there is good evidence that explicit intrinsic information can be extracted from our latent intrinsics. Our method clearly “knows” albedo, and this information can be elicited without examples. Similarly, it “knows” something about luminaires, such as their locations and effects. It is intriguing to speculate that it “knows” other information relevant to relighting, such as depth or surface microstructure. Future work will pursue this line of inquiry and also focus on developing a purely unsupervised framework to infer intrinsic and extrinsic properties from collections of in-the-wild images. This will include refining probing techniques for better extraction of explicit intrinsics and identifying additional intrinsic properties crucial for relighting that do not align with the current taxonomy. We believe this will improve the applicability and robustness of our approach, making it suitable for a wider range of real-world scenarios.
@@ Acknowledgment
AB thanks Stephan R. Richter for the discussions that led to the consideration of intrinsic images as latent variables. This material in part is based upon work supported by the National Science Foundation under Grant No. 2106825, and by a gift from Boeing.